# ArtNVG: Content-Style Separated Artistic Neighboring-View Gaussian Stylization

### v1

@@ 3.2.1 Overview
Our method starts with rendering content images from the original 3DGS scene, followed by the Content-Style Separated Control, which encodes the content images and style image separately using the projection modules from CSGO [47] and a Tile ControlNet[56]. Simultaneously, we sample and group neighboring views to ensure local consistency. During the denoising process, the down blocks of the UNet mainly control the content, whereas the up blocks control the style. Thus, the content and style controls are passed to different cross-attention layers. As for information among neighboring views, it is shared in self-attention layers. Finally, stylized images output by the neighboring-view diffusion model are used to finetune the 3D scene and we apply NNFM loss proposed by Artistic Radiance Fields (ARF) [55] to enhance style alignment and details in stylization.
@@ 3.2.2 Content-Style Separated Control
To stylize a 3D scene, we need to ensure effective style transfer and accurate content preservation. In 3D stylization tasks, one of the main concerns for everyone is the risk of content information leakage of the style image and style information leakage of content images. To reduce the risk of unwanted information leakage, we encode the style image and content images separately and inject controls through cross-attention layers.
Style Control Module. Our method uses the CSGO style projection module as the style control module. The style control can be denoted as $\mathcal{D}_{s}=\mathcal{F}_{s}(\mathcal{E}(I_{\mathrm{sty}}))$, where $I_{\mathrm{sty}}$ is the style image, $\mathcal{E}$ is the image encoder, and $\mathcal{F}_{s}$ is the CSGO style projection module.
@@ 3.2.3 Attention-based Neighboring-View Alignment
While our content control module offers effective content control, it is still difficult to ensure alignment among neighboring views. Despite sharing the same style control and similar content controls, stylized images nearby may exhibit discrepancies in details, causing locally inconsistent colors and textures in scenes. Inspired by previous studies that achieve view consistency by manipulating attention [24, 46, 11], we propose an Attention-based Neighboring-View (NV) Alignment mechanism by modifying self-attention layers into Neighboring-View attention layers in the diffusion model. With attention sharing, images are no longer stylized alone. Instead, images of neighboring views are jointly stylized as a group.
@@ 3.3 Implementation Details
Our method is implemented by using the PyTorch library. For 3D reconstruction and visualization, we use models and tools from the NeRFStudio [39] library. Specifically, we employ the ”splatfacto” model for 3D reconstruction. Following settings in CSGO [47], we employ Stable Diffusion XL v1.0 and pre-trained weight of CSGO models for 2D image stylization using the Diffusers [41] library.
@@ 4.1 Experiment Setup
Baselines. We compare our method with the state-of-the-art 3DGS stylization method StyleGaussian [23]. StyleGaussian is a feedforward-based framework, which embeds 3DGS and does stylization on the embeddings. We use its pre-trained models for evaluation.
@@ 4.2 Qualitative Results
We demonstrate the qualitative results in Fig. 4, showcasing the advanced capabilities of ArtNVG in achieving superior style transfer performance. ArtNVG excels in creating high-quality 3D scenes with enhanced alignment to the artistic style of reference images while preserving high content fidelity. StyleGaussian struggles to prevent information leakage in rows 3 and 4, which is likely to originate from the style transfer process of embedded Gaussians. In addition, StyleGaussian fails to generate 3D scenes with high quality, resulting in artifacts such as too many floaters (shown in row 1), blurring details (shown in rows 2 and 6), and locally inconsistent colors (shown in rows 3, 4, 5 and 6). This phenomenon may be largely due to StyleGaussian’s decoded stylized RGB design, by which the quality of the generated scenes depends on the decoded features. Its features are extracted by VGG, resulting in the limitation of feature representation. In contrast, our method uses Content-Style Separated Control to avoid information leakage and incorporates Attention-based Neighboring-View Alignment to promise locally consistent colors and textures, leading to fewer artifacts in scenes.
@@ 4.3 User Study
Figure 3: User Study. We record the user preference of our method and the baseline. Our method obtains more preference in content fidelity, style alignment, and visual quality than the baseline.
We conducted a user study comparing our method with StyleGaussian. We randomly chose 10 sets of stylized views rendered by different methods and invited 20 participants. For each view, there are three evaluation indicators: content fidelity, style alignment and visual quality. Then, participants were asked to vote for a view based on these indicators. We collected a total of 200 votes for each indicator and demonstrated the results in Fig. 3. Our method outperforms StyleGaussian across all evaluation indicators, indicating our method can perform excellent artistic style transfer with high visual fidelity.
Figure 4: Qualitative Results of Comparison. We show diverse results of stylization in various scenes with different styles. Compared with StyleGaussian [23], our method generates locally consistent scenes with better style alignment while preserving high content fidelity.
@@ 4.4 Quantitative Results
Table 1: Quantitative Results of Comparison. We evaluate the performance of our method against StyleGaussian [23] in terms of content fidelity, style similarity and temporal consistency, using CFSD ($\downarrow$), CSD score ($\uparrow$) and CLIP-DC ($\uparrow$).
While the nature of stylization is inherently subjective, we mostly rely on the qualitative results and user study for evaluation. Nevertheless, motivated by [11, 13], we apply auxiliary quantitative metrics over 8 scenes with different styles to evaluate our method more comprehensively. The quantitative results are summarized in Tab. 1. Our method outperforms StyleGaussian on all evaluation metrics. Especially for content fidelity, CFSD of our method is much lower than StyleGaussian, which means our method has a better capability of capturing content features. This may be largely due to the powerful content control module used in our method, whereas StyleGaussian only uses VGG features. For style alignment and multi-view consistency, our method also achieves higher CSD score and CLIP-DC than StyleGaussian.
@@ 4.5 Ablation Study
w/o NV Attention. To investigate the importance of maintaining the local consistency of stylized images prior, we conducted an ablation by disabling Neighboring-View attention. As evidenced by in Fig. 5 (c), this variant harms the visual quality, leading to blurry details in finetuned 3DGS scenes, which is largely due to locally inconsistent details in neighboring views. It can also be found that there is a decline in all metrics in Tab. 2.
@@ 6 Conclusion
This document serves as supplementary materials for ArtNVG: Content-Style Separated Artistic Neighboring-View Gaussian Stylization, providing additional implementation details, metrics, and extended results. We highly recommend that reviewers view the attached videos to fully assess the performance of our stylization method across multiple perspectives.

### v2

@@ 3.2.1 Overview
Our method starts with rendering content images from the original 3DGS scene, followed by the Content-Style Separated Control, which encodes the content images and style image separately using the projection modules from CSGO [47] and a Tile ControlNet [56]. Simultaneously, we sample and group neighboring views to ensure local consistency. During the denoising process, the down blocks of the UNet mainly control the content, whereas the up blocks control the style. Thus, the content and style controls are passed to different cross-attention layers. As for information among neighboring views, it is shared in self-attention layers in both the UNet and the Tile ControlNet, so we replace the original self-attention layers with the Neighboring-View Attention layers to maintain neighboring view consistency. Finally, stylized images output by the neighboring-view diffusion model are used to finetune the original 3D scene and we apply NNFM loss proposed by Artistic Radiance Fields (ARF) [55] to enhance style alignment and details in stylization.
@@ 3.2.2 Content-Style Separated Control
To stylize a 3D scene, we need to ensure effective style transfer and accurate content preservation. In 3D stylization tasks, one of the main concerns for everyone is the risk of content information leakage of the style image and style information leakage of content images. Previous methods [11, 52] use an inversion-based content preservation approach, which adds noises to the latents of original renderings as the input of the diffusion model, posing a high risk of content information leakage. To reduce the risk of unwanted information leakage, we encode the style image and content images separately and inject controls through cross-attention layers.
Style Control Module. Our method uses the CSGO style projection module as the style control module. The style control can be denoted as $\mathcal{D}_{s}=\mathcal{F}_{s}(\mathcal{E}(I_{\mathrm{sty}}))$, where $I_{\mathrm{sty}}$ is the style image, $\mathcal{E}$ is the CLIP image encoder, and $\mathcal{F}_{s}$ is the CSGO style projection module.
@@ 3.2.3 Attention-based Neighboring-View Alignment
While our content control module offers effective content control, it is still difficult to ensure alignment among neighboring views. Neighboring views refer to the views, of which camera centers are closer, meaning the corresponding images have larger overlapped regions. Although neighboring views share identical style controls and similar content controls due to substantial region overlapping, their stylized outputs may demonstrate noticeable discrepancies among them. This phenomenon triggers arbitrary fine-tuning in overlapping areas, resulting in color and texture discontinuities in detail across the scene. Inspired by previous studies that achieve view consistency by manipulating attention [24, 46, 11], we propose an Attention-based Neighboring-View (NV) Alignment mechanism by modifying self-attention layers into Neighboring-View attention layers in the diffusion model. Our approach demonstrates superior neighboring view consistency compared to the aforementioned approaches, as it employs attention sharing mechanism based on camera center distance rather than relying on random selection or predefined manual selection. With attention sharing, images are no longer stylized alone. Instead, images of neighboring views are jointly stylized as a group and as a result, details are more consistent among neighboring views (demonstrated in Fig. 3).
Figure 3: Comparative Analysis of Detail Consistency in Neighboring Views with and without Neighboring-View Attention. This illustration presents a comparative visualization of stylization results between two neighboring views. The regions demarcated by red bounding boxes clearly exhibit enhanced detail consistency when NV Attention is implemented, in contrast to the results obtained without NV Attention.
@@ 3.3 Implementation Details
Our method is implemented by using the PyTorch library. For 3D reconstruction and visualization, we use models and tools from the NeRFStudio [39] library. Specifically, we employ the “splatfacto” model for 3D reconstruction. Following settings in CSGO [47], we employ Stable Diffusion XL v1.0 and pre-trained weight of CSGO models for 2D image stylization using the Diffusers [41] library.
@@ 4.1 Experiment Setup
Baselines. We compare our method with the state-of-the-art 3DGS stylization methods StyleGaussian [23] and InstantStyleGaussian [52]. StyleGaussian is a feedforward-based framework, which embeds 3DGS and does stylization on the embeddings. We use its pre-trained models for evaluation. InstantStyleGaussian is also based on diffusion stylization, which can serve as an effective comparative baseline for evaluation. We implement this method through a faithful reproduction of its original framework, adhering to its original specifications in its publication.
@@ 4.2 Qualitative Results
Figure 4: Qualitative Results of Comparison. We show diverse results of stylization in various scenes with different styles. Compared to SOTA methods StyleGaussian and InstantStyleGaussian, our method achieves the best visual quality by generating locally consistent scenes with better style alignment while preserving high content fidelity.
We demonstrate the qualitative results in Fig. 4, showcasing the advanced capabilities of ArtNVG in achieving superior style transfer performance. ArtNVG excels in creating high-quality 3D scenes with enhanced alignment to the artistic style of reference images while preserving high content fidelity. StyleGaussian struggles to prevent information leakage in rows 3 and 4, which is likely to originate from the style transfer process of embedded Gaussians. In addition, StyleGaussian fails to generate 3D scenes with high quality, resulting in artifacts such as too many floaters (shown in row 1), blurring details (shown in rows 2 and 6), and locally inconsistent colors (shown in rows 3, 4, 5 and 6). This phenomenon may be largely due to StyleGaussian’s decoded stylized RGB design, by which the quality of the generated scenes depends on the decoded features. Its features are extracted by VGG, resulting in the limitation of feature representation. InstantStyleGaussian performs better than StyleGaussian in content preservation and visual quality. However, it still suffers from limitations like blurring details (shown in rows 1, 2 and 4), severe content leakage (shown in row 4), and even occasional failure to reconstruct a meaningul 3D scene (shown in row 6), which may originates from its neglect of neighboring view consistency and adoption of an inversion-based content preservation approach in diffusion stylization. It preserves the content information by adding noises on the original renderings following the implementation of some previous 3D editing frameworks [13, 8, 40, 46], but this way of content preservation inherently poses a high risk of leaking the style information of content images. Besides, it does not take the neighboring view consistency into consideration, causing locally arbitrary details in stylized images. In contrast, our method uses Content-Style Separated Control to avoid information leakage and incorporates Attention-based Neighboring-View Alignment to promise locally consistent colors and textures, leading to fewer artifacts in scenes.
@@ 4.3 User Study
Figure 5: User Study. We record the user preference of our method and the baselines. Our method obtains more preference in content fidelity, style alignment, and visual quality than the baselines.
We conducted a user study comparing our method against StyleGaussian and InstantStyleGaussian. We randomly chose 10 sets of stylized views rendered by different methods and invited 20 participants. For each view, there are three evaluation indicators: content fidelity, style alignment and visual quality. Then, participants were asked to vote for a view based on these indicators. We collected a total of 200 votes for each indicator and demonstrated the results in Fig. 5. Our method outperforms StyleGaussian and InstantStyleGaussian across all evaluation indicators, indicating our method can perform excellent artistic style transfer with high visual fidelity.
@@ 4.4 Quantitative Results
Table 1: Quantitative Results of Comparison. We evaluate the performance of our method against StyleGaussian and InstantStyleGaussian in terms of content fidelity, style similarity and temporal consistency, using CFSD ($\downarrow$), CSD score ($\uparrow$) and CLIP-DC ($\uparrow$). Our method surpasses them on all the metrics.
While the nature of stylization is inherently subjective, we mostly rely on the qualitative results and user study for evaluation. Nevertheless, motivated by [11, 13], we apply auxiliary quantitative metrics over 8 scenes with different styles to evaluate our method more comprehensively. The quantitative results are summarized in Tab. 1. Our method outperforms StyleGaussian and InstantStyleGaussian on all evaluation metrics. Particularly in content fidelity, our approach achieves significantly lower CFSD scores, compared to both StyleGaussian and InstantStyleGaussian, indicating enhanced capability in preserving structural content features. Regarding style alignment and multi-view consistency, our method obtains substantially higher CSD and CLIP-DC scores, surpassing the baselines. These improvements primarily stem from the implementation of the Neighboring-View Attention mechanism, which effectively maintains cross-view consistency, and the integration of the CSGO model for enhanced feature representation. In contrast, StyleGaussian relies solely on VGG features, while InstantStyleGaussian is limited to an inversion-based InstantStyle model, both of which lack effective design for maintaining view consistency, content fidelity and style alignment.
@@ 4.5 Ablation Study
w/o NV Attention. To investigate the importance of maintaining the local consistency of stylized images, we conducted an ablation by disabling Neighboring-View attention. As evidenced by in Fig. 6 (c), this variant harms the visual quality, leading to blurry details in finetuned 3DGS scenes, which is largely due to locally inconsistent details in neighboring views. It can also be found that there is a decline in all metrics in Tab. 2.
@@ 6 Conclusion
This document serves as supplementary materials for ArtNVG: Content-Style Separated Artistic Neighboring-View Gaussian Stylization, providing additional implementation details, metrics, and extended results. We highly recommend that reviewers view the following results in App. C to fully assess the performance of our stylization method across multiple perspectives.